OpenStudy (anonymous):

A business owner is looking at sales records for the last 10 years. In 2009, he calculated his net profit was $75,000 for the year. In 1999, he did not make a profit, instead making a net loss of $5,000. Find the rate of change of his profit from 1999 to 2009. A. $35/yr B. $40/yr C. $7,000/yr D. $8,000/yr

OpenStudy (ranga):

In 2009, profit = +$75,000 In 1999, profit = -$5,000 (loss can be written as negative profit) The change in the profit is: $75,000 - (-$5,000) = 75,000 + 5,000 = 80,000 This happens over a 10 year period (2009 - 1999 = 10) Rate of change = change in profit / number of years = 80,000 / 10 = ?

OpenStudy (anonymous):

2009 profit of $75,000 - l999 loss of $5,000 = $70,000 profit past 10 years $70,000/10 years = $7000 per year profit.

OpenStudy (anonymous):

is the answer D @ranga

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

you are welcome.